OpenStudy (anonymous):

Evaluate 12 - 6 + (14 - 8)2 = ____________.

OpenStudy (anonymous):

18

OpenStudy (anonymous):

the last two is an exponent. can you help? @prototype6914

OpenStudy (dejmccray):

what do u need help with ?

OpenStudy (anonymous):

answering it I dont understand. @dejmccray

OpenStudy (dejmccray):

ok so 12-6+(14-8)2= first do 14-8 which is 6 so now its 12-6+6*2 now do multiplication which would be 6*2=12 now its 12-6+12 now just subtract and add 12-6=6+12=18 answer is 18